### v1

# FuzzySQL: Uncovering Hidden Vulnerabilities in DBMS Special Features with LLM-Driven Fuzzing

Abstract.
Traditional database fuzzing techniques primarily focus on syntactic correctness and general SQL structures, leaving critical yet obscure DBMS features, such as system-level modes (e.g., GTID), programmatic constructs (e.g., PROCEDURE), advanced process commands (e.g., KILL), largely underexplored. Although rarely triggered by typical inputs, these features can lead to severe crashes or security issues when executed under edge-case conditions. In this paper, we present FuzzySQL, a novel LLM-powered adaptive fuzzing framework designed to uncover subtle vulnerabilities in DBMS special features. FuzzySQL combines grammar-guided SQL generation with logic-shifting progressive mutation, a novel technique that explores alternative control paths by negating conditions and restructuring execution logic, synthesizing structurally and semantically diverse test cases. To further ensure deeper execution coverage of the back end, FuzzySQL employs a hybrid error repair pipeline that unifies rule-based patching with LLM-driven semantic repair, enabling automatic correction of syntactic and context-sensitive failures. We evaluate FuzzySQL across multiple DBMSs, including MySQL, MariaDB, SQLite, PostgreSQL and Clickhouse, uncovering 37 vulnerabilities, 7 of which are tied to under-tested DBMS special features. As of this writing, 29 cases have been confirmed with 9 assigned CVE identifiers, 14 already fixed by vendors, and additional vulnerabilities scheduled to be patched in upcoming releases. Our results highlight the limitations of conventional fuzzers in semantic feature coverage and demonstrate the potential of LLM-based fuzzing to discover deeply hidden bugs in complex database systems.
Database security; Fuzzing; AI for security; Language validity
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacy Database and storage securityccs: Security and privacy Software and application security
1. Introduction
Modern database management systems (DBMSs) sit at the core of software infrastructure that is critical to security and reliability, where subtle logic flaws can lead to persistent crashes, state corruption, and severe service disruption. As a result, fuzzing has emerged as an effective technique for uncovering DBMS bugs by generating SQL statements and executing them against target systems. However, most existing DBMS fuzzers (37; R. Zhong, Y. Chen, H. Hu, H. Zhang, W. Lee, and D. Wu (2020); M. Rigger and Z. Su (2020c); Y. Liang, S. Liu, and H. Hu (2022); Z. Jiang and Z. Su (2024)) predominantly focus on general-purpose statement structures (e.g., SELECT), covering only a limited portion of the DBMS grammar. Recent efforts such as BUZZBEE (Yang et al., 2024) and POLYGLOT (Chen et al., 2021) emphasize cross-target generalizability, rather than systematically exercising DBMS-specific features such as read-only modes or GTID control.
DBMS special features are vendor-specific extensions that deviate from standard SQL specifications and, although invoked less frequently, are deeply embedded in core execution paths and can trigger severe system failures under edge conditions. For example, Listing 1 shows that a seemingly benign combination of statements, such as UPDATE HISTOGRAM and RESET within a read-only transaction, can lead to a full-system crash—behaviors that are rarely exercised by conventional fuzzers. Existing DBMS fuzzers have explored SQL generation and mutation from multiple perspectives, including structural complexity (Zhong et al., 2020; Jiang et al., 2023), sequence-level interactions (Fu et al., 2022; Liang et al., 2023), and oracle-guided logic testing (Rigger and Su, 2020a; Liang et al., 2022; Jiang and Su, 2024). Despite these advances, they share a common limitation: none have made effective attempts to systematically explore feature-specific logic. At a fundamental level, this limitation stems from the mismatch between feature-specific DBMS behaviors and the design assumptions of existing fuzzers, and manifests in several aspects.
Concretely, we identify three key factors that hinder existing fuzzers from effectively exploring feature-specific DBMS logic. First, traditional fuzzers lack support for feature-specific SQL constructs that are often tied to internal DBMS mechanisms. These constructs are not only vendor-specific but also version-dependent, making rule-based modeling costly and slow to adapt. Second, semantic modeling remains a fundamental challenge. Without understanding complex schema dependencies and execution context, traditional tools struggle to synthesize auxiliary objects (e.g., PROCEDURE) or establish the preconditions needed to exercise deep execution paths, leading to low coverage of feature-specific logic. Third, existing fuzzers typically lack global state tracking and replay mechanisms, making it difficult to identify compound failures that arise only through interactions across multiple statements or test cases. For example, the underlying bug demonstrated by Listing 1 was triggered by a latent interaction among statements from four distinct test cases.
To address these limitations, we present FuzzySQL, a novel adaptive fuzzing framework driven by LLMs and designed to systematically uncover vulnerabilities in the under-tested feature space of modern DBMSs. FuzzySQL introduces the following key innova-
1CREATE TABLE t0 (xxx);
2ALTER DATABASE test_db READ ONLY = 1;
3ANALYZE TABLE t0 UPDATE HISTOGRAM ON v1;
4RESET BINARY LOGS AND GTIDS;
5INSERT INTO t0 VALUES (xxx);
Listing 1: Motivational example of a MySQL vulnerability related to DBMS special features.
tions. First, FuzzySQL targets feature-specific SQL constructs by combining grammar-guided template expansion with LLM-powered instantiation (Section 3.1), enabling systematic coverage of vendor-specific and version-dependent DBMS features that are difficult to model using fixed rules. Second, to overcome the challenges of semantic modeling, FuzzySQL performs semantic exploration through a unified design that integrates progressive mutation with automated error repair (Sections 3.2 and 3.3). Specifically, it prioritizes semantic diversity by recombining logic fragments across seed programs and applying conditional logic shifts, while automatically repairing high-value test cases that fail due to missing objects or unmet semantic preconditions. This design enables FuzzySQL to construct semantically valid and diverse SQL sequences that reach deep, feature-dependent execution paths. Finally, FuzzySQL maintains the global execution context throughout fuzzing and performs replay-based analysis to identify and isolate compound failures that arise only under specific internal DBMS states (Section 3.4). This paper makes the following key contributions:
  • Uncovering a critical blind spot in DBMS fuzzing. We demonstrate that conventional fuzzers typically neglect the special features of DBMSs, such as GTID handling, which can lead to severe failures, but are rarely triggered by general-purpose test cases.
  • Enabling deep semantic exploration through LLM-driven fuzzing. We present FuzzySQL, an adaptive fuzzing framework that combines template-guided generation, context-aware mutation, and a hybrid error-repair pipeline. This design enables the synthesis of structurally valid and semantically rich SQL statement sequences that activate deeper DBMS control paths.
  • Comprehensive feature coverage via execution state maintenance and replay. FuzzySQL flexibly adapts to diverse DBMS dialects and feature sets and maintains the global execution context during fuzzing and crash validation. This allows it to detect feature-related bugs that depend on internal state transitions and diverse execution environments.
  • Discovering real-world vulnerabilities across multiple DBMSs. We evaluate FuzzySQL on MySQL, MariaDB, SQLite, and ClickHouse, uncovering 37 vulnerabilities, of which 29 have been confirmed by developers and 7 are associated with under-tested DBMS features. These results highlight FuzzySQL’s ability to expose semantic and feature-specific vulnerabilities consistently missed by existing approaches.
2. Motivation
Modern DBMSs implement a wide range of system-specific features that go far beyond standard SQL support, introducing complex internal mechanisms and stateful behaviors. These under-tested features often expose latent vulnerabilities that are difficult to uncover through traditional fuzzing techniques. In this section, we first highlight the limitations of existing fuzzers in exercising such features through a motivating example. We then discuss why previous approaches fail to trigger these bugs and explore the potential of LLMs to address the associated challenges.
2.1. Under-Tested Features in Modern DBMSs
While traditional database fuzzers primarily target general-purpose SQL statements such as SELECT, INSERT, and CREATE TABLE, modern DBMSs expose a broader attack surface through system-level features that manipulate internal state and metadata. Listing 1 presents a minimal SQL sequence that triggers a persistent crash in MySQL by misusing such special-purpose statements.
These features include operations for toggling system modes (e.g., READ ONLY), managing replication state (e.g., BINARY LOGS and GTID), and maintaining internal statistics (e.g., UPDATE HISTOGRAM). Although invoked less frequently than standard DML and DDL statements, they directly interact with core DBMS subsystems and global execution state. Each statement in Listing 1 is syntactically valid and semantically reasonable in isolation. However, their specific combination exposes a subtle control-path vulnerability in MySQL’s metadata lock management. In particular, toggling READ ONLY mode and updating table histograms interact with metadata locks maintained by the dictionary subsystem, while resetting binary logs alters session-level replication state. When followed by a regular INSERT, these interactions violate internal lock invariants and trigger an assertion failure in the metadata lock checker.
More critically, triggering this vulnerability leads to persistent state corruption that survives across subsequent executions. Even after restarting the server, later queries result in immediate crashes, indicating a lasting impact on internal state rather than a transient fault. This example demonstrates that under-tested DBMS features can interact in unexpected ways, producing severe and state-dependent failures that are difficult to expose through traditional statement-level fuzzing.
2.2. Why Existing Fuzzers Miss These Bugs
Despite progress in DBMS fuzzing, most existing tools remain centered on crafting syntactically valid DML and DDL statements, with limited awareness of feature-specific SQL constructs. Moreover, such features vary significantly across DBMSs and evolve rapidly between releases. For example, RESET BINARY LOGS AND GTIDS in Listing 1 is a recent feature introduced in MySQL 8.4 to replace the legacy RESET MASTER, making manual rule-based support both costly and prone to lag behind DBMS evolution.
Figure 1. Overview of FuzzySQL.
In addition, prior tools typically lack global state awareness. Without tracking execution context or supporting replay, they may overlook or misclassify state-dependent failures as false positives. In our investigation, the vulnerability shown in Listing 1 could not be triggered by a single test case; instead, it was reconstructed from four independent inputs identified among thousands of PoCs. These inputs were discovered through sequential replay and input minimization, as there was no semantic guidance suggesting that READ ONLY toggles and RESET commands should be combined. This limitation underscores that current fuzzers are fundamentally inadequate for exposing deep, feature-related vulnerabilities.
2.3. LLM: Opportunities and Challenges
LLMs enable semantic understanding of complex SQL constructs, allowing them to generate coherent multi-statement SQL sequences, infer implicit dependencies, and synthesize missing objects (e.g., creating a PROCEDURE when a CALL is issued). This reduces reliance on manually crafted rules in DBMS fuzzing and supports more comprehensive testing of system-specific features. However, LLMs also introduce unique challenges in the fuzzing context. Due to their exposure to well-formed data during pretraining and instruction tuning, they exhibit a strong bias toward generating syntactically correct, semantically plausible, and conventionally safe inputs. Specific design choices are required to guide LLMs toward semantic reasoning over schema context, rather than defaulting to generic low-risk outputs.
A notable characteristic of LLMs is hallucination, where models generate syntactically or semantically invalid content relative to the underlying system or schema. While typically undesirable, hallucinations can be beneficial for fuzzing, as they help explore unexpected behaviors and may expose corner-case bugs. In the fuzzing context, hallucination acts as a feature that promotes exploration beyond normative inputs. We further illustrate this phenomenon in Section 5.4.
2.4. Goals of FuzzySQL
Based on the motivations described above, we design FuzzySQL as an LLM-powered adaptive fuzzing framework that aims to cover the under-tested feature space of DBMS. FuzzySQL is designed to:
  1. i.
    Flexibly adapt to diverse DBMS dialects and system-specific features through grammar-based template expansion and LLM-guided semantic instantiation.
  2. ii.
    Generate syntactically valid and semantically meaningful SQL sequences that invoke diverse internal behaviors.
  3. iii.
    Maintain global execution context and support replay-based PoC validation to detect and isolate persistent, state-dependent bugs rooted in non-trivial control flows.
FuzzySQL aims to shift the paradigm of database fuzzing from traditional syntax-driven or random mutation strategies toward a semantically guided approach that emphasizes deeper program logic and broader exploration of system features. We provide an anonymized artifact repository to support reproducibility, as described in the Open Science Appendix C.
3. Methodology
To uncover deep and stateful bugs in modern DBMSs, FuzzySQL adopts a unified fuzzing pipeline that integrates structured generation, semantic mutation, automated repair, and replay-based validation, as illustrated in Figure 1. Rather than generating isolated SQL statements, FuzzySQL continuously evolves multi-statement programs under a persistent execution context, enabling exploration of feature-specific and state-dependent behaviors. Fuzzing begins with grammar-guided SQL template expansion, followed by LLM-based instantiation under the current schema state to produce executable and semantically meaningful inputs. Executed test cases that yield new coverage are retained and further evolved via progressive mutation, which recombines statement sequences and applies logic-level transformations to explore alternative control paths. When execution fails due to syntactic or semantic inconsistencies, FuzzySQL applies a multi-stage repair pipeline to salvage high-value inputs instead of discarding them. Finally, detected crashes are validated through replay-guided execution and minimized to concise, reproducible PoCs that capture state-dependent failures.
3.1. Grammar-Guided SQL Generation
To generate diverse and structurally valid SQL inputs, FuzzySQL employs a two-phase grammar-driven generation process. First, it performs a recursive expansion of production rules extracted from the DBMS grammar definitions. Second, it uses LLM to instantiate concrete SQL statements guided by both the expanded templates and the contextual schema state.
Figure 2. CFG rules of the ‘alterTable’ statement.
3.1.1. Grammar Template Expansion
To systematically cover the SQL grammar space, FuzzySQL first parses the target DBMS’s ANTLR4 grammar specification to extract production rules. We abstract all rules into four categories to enable controlled and interpretable expansion, as illustrated in Figure 2:
  • SequenceRule: A sequence of grammar symbols to be expanded in order.
  • OptionalRule: Grammar elements that may be omitted or included (e.g., IGNORE).
  • ChoiceRule: Mutually exclusive options, where one is selected from a set (e.g., ONLINE vs. OFFLINE).
  • RepeatRule: Patterns that can be repeated one or more times (e.g., comma-separated lists of alterSpecification).
1alterTable -> ALTER ONLINE TABLE [tableName] [alterSpecification] [partitionDefinitions]
2alterTable -> ALTER TABLE [tableName] [alter-Specification],[alterSpecification]
Listing 2: SQL templates derived from the alterTable rule.
Based on this classification, we scan and normalize the grammar file, enabling heuristic control during recursive expansion. Then, FuzzySQL constructs recursive derivation trees rooted at top-level statement nonterminals. Two safeguards are applied to ensure diversity and manageability: (i) A configurable maximum recursion depth to bound nested grammar derivations (ii) Rule-specific usage quotas to prevent structural redundancy (e.g., excessive parentheses or nested subqueries). The final output is a collection of skeletal SQL templates—syntax-valid but uninstantiated statement blueprints. Listing 2 shows examples of grammar templates expanded from the alterTable rule. Both templates can be further expanded through the alterSpecification rule, as long as the depth and quota limits are not exceeded. Grammar templates form the structural backbone of FuzzySQL’s seed generation process, ready to be instantiated with real schema and data context by LLMs in the next phase.
3.1.2. LLM-Powered Fuzzy Instantiation
Given a grammar-expanded template, FuzzySQL uses a LLM to instantiate the abstract statement into a complete executable SQL query. To guide this process, we maintain a dynamic schema context—populated via randomly generated CREATE TABLE and INSERT statements, which includes table names, column types and relational structures. These initialization statements not only establish a valid database environment with rich type coverage (e.g., INT, ENUM, GEOMETRY), but also provide the semantic context necessary for the LLM to generate coherent and meaningful SQL logic.
For example, when instantiating a template involving a JOIN or subquery, the LLM leverages both the schema and the grammatical scaffold to synthesize a valid and meaningful test case. Unlike traditional fuzzers, which randomly inject literals or select columns uniformly, FuzzySQL is context-aware and schema-consistent. As a result, FuzzySQL can generate a wide variety of semantically correct and structurally diverse SQL queries ranging from simple data manipulation to complex administrative commands while ensuring that they are executable in the current DBMS state.
Algorithm 1 Logic-Shifting Progressive Mutation
1:Input: Two SQL test cases $\mathcal{T}_{1}$, $\mathcal{T}_{2}$
2:Output: Mutated test case $\mathcal{T}^{*}$
3: Extract schema fragments $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ and operations $\mathcal{O}_{1}$, $\mathcal{O}_{2}$
4: Initialize empty schema set $\mathcal{S}\leftarrow\emptyset$
5:for each table $t$ in $\mathcal{S}_{1}\cup\mathcal{S}_{2}$ do
6:  if $t$ exists in both $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ then
7:   Randomly select one version of $t$ from $\mathcal{S}_{1}$ or $\mathcal{S}_{2}$
8:  else
9:   Use the available definition of $t$
10:  end if
11:  Append chosen schema statements of $t$ into $\mathcal{S}$
12:end for
13: Initialize empty operation sequence $\mathcal{O}\leftarrow EmptyList$
14: Let $i,j$ be indices for $\mathcal{O}_{1}$, $\mathcal{O}_{2}$
15:while $i<|\mathcal{O}_{1}|$ or $j<|\mathcal{O}_{2}|$ do
16:  if $i<|\mathcal{O}_{1}|$ and ($j\geq|\mathcal{O}_{2}|$ or random <0.5<0.5) then
17:   Append $\mathcal{O}_{1}[i]$ to $\mathcal{O}$; $i\leftarrow i+1$
18:  else
19:   Append $\mathcal{O}_{2}[j]$ to $\mathcal{O}$; $j\leftarrow j+1$
20:  end if
21:end while
22:for each statement $s$ in $\mathcal{O}$ do
23:  Sample a drop probability $p_{\text{drop}}\sim\mathcal{U}(0.2,0.4)$
24:  if $random()<p_{\text{drop}}$ then
25:   Drop $s$
26:  end if
27:end for
28:for each statement $s$ in $\mathcal{O}$ do
29:  Apply conditional rewrites $(k\rightarrow k^{\prime})$:
30:   if $k$ is JOIN, randomly replace with supported JOIN;
31:   else replace with $k^{\prime}$
32:end for
33:Return: $\mathcal{T}^{*}=\mathcal{S}$ $\cup$ $\mathcal{O}$
3.2. Logic-Shifting Progressive Mutation
While grammar-based generation ensures structural diversity, mutation plays a crucial role in evolving test cases toward deeper execution paths. FuzzySQL introduces context reconstruction that operates at the SQL sequence level, incorporating a logic-shifting progressive mutation strategy to maximize semantic diversity and minimize redundancy. This design avoids complex parsing and rigid mutation rules, enabling efficient discovery of diverse behaviors across DBMS dialects with minimal seed input requirements. The overall process is illustrated in Algorithm 1.
3.2.1. Sequence Crossover Synthesis
To generate new test cases, FuzzySQL randomly selects a pair of seed SQL statement sequences from the seed pool and performs schema-aware crossover synthesis. The process follows 3 steps: Schema Unification. As shown in lines 3–11 of Algorithm 1, we first extract schema initialization statements (e.g., CREATE TABLE, INSERT INTO) from both $T_{1}$ and $T_{2}$, denoted as $S_{1}$ and $S_{2}$. For each table identifier $t$ appearing in $S_{1}\cup S_{2}$, we resolve potential conflicts: if $t$ is defined in both seeds, one version is randomly chosen to ensure consistency and eliminate redundancy. The unified schema statements are appended to form a consistent schema set $S$.
Sequence Crossover. We extract non-initialization SQL statements (e.g., SELECT, ANALYZE, KILL) from $T_{1}$ and $T_{2}$ to form operation sequences $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$. As shown in Lines 13–21 of Algorithm 1, these sequences are probabilistically interleaved to produce a new sequence $\mathcal{O}$ that preserves partial order while enabling cross-seed logic recombination. This allows FuzzySQL to synthesize test cases that blend behaviorally diverse fragments and explore context-dependent interactions. Sequence crossover may introduce semantic inconsistencies, such as mismatched table or column references. To reduce identifier conflicts, we adopt a unified naming convention during generation (e.g., t, c with numeric suffixes). Moreover, syntactic or semantic errors introduced during mutation are not fatal—they are addressed in the subsequent error repair stage.
Size and Diversity Control. To avoid overly large or redundant test cases, we perform probabilistic dropping of operations (lines 22–27). For each statement $s$ in $\mathcal{O}$, a drop probability $p_{\text{drop}}$ is sampled from a uniform distribution $\mathcal{U}$, and $s$ is discarded with that probability. This strategy ensures that the generated test cases remain compact, computationally efficient, and support progressive exploration of novel context combinations. The resulting synthetic case blends the execution contexts and logic flows from both parents, increasing the likelihood of triggering corner-case bugs that require complex state transitions.
3.2.2. Conditional Logic Shift
After crossover synthesis, FuzzySQL applies a second-stage mutation phase that targets the logical structure of SQL conditions. This logic-shifting transformation rewrites common predicates and joins patterns to explore alternative semantic paths while preserving syntactic validity. The process is implemented via rule-driven string rewrites, requiring no schema-specific reasoning or AST instrumentation, and generalizes well across SQL dialects. This logic-level rewriting corresponds to lines 28–32 in Algorithm 1, where each statement $s$ in the filtered operation sequence is examined for conditional rewrite opportunities. Specifically, FuzzySQL maintains a rewrite rule set $\{k\rightarrow k^{\prime}\}$, and applies them conditionally based on statement type. We organize these transformations into two primary categories:
  1. i.
    Predicate mutations, which rewrite conditional operators (e.g., = vs. !=, IN vs. NOT IN, IS NULL vs. IS NOT NULL, or ordering directions) in WHERE, HAVING, and ORDER BY clauses. These lightweight changes preserve syntactic structure while enabling exploration of complementary execution paths.
  2. ii.
    JOIN rewrites, which substitute join variants (e.g., INNER, LEFT, or CROSS JOIN) to perturb query structure and query-plan generation, potentially exposing logic differences in optimization or error-handling stages.
In contrast to AST-based fuzzers such as Squirrel (Zhong et al., 2020) that rely on heavyweight intermediate representations, our mutation layer is minimal and DBMS-agnostic, allowing efficient fuzzing over LLM-generated SQL with low engineering cost. To maintain cross-system compatibility, FuzzySQL includes dialect-aware filtering: for instance, disabling FULL JOIN in SQLite or adapting join syntax for ClickHouse. This ensures that semantic variation remains within valid syntactic boundaries for each target engine.
3.3. Automated Error Repair
Instantiating complex grammar templates into fully valid SQL statements in a single pass is inherently difficult, particularly for under-tested features such as PROCEDURE. In addition, random mutations often introduce contextual inconsistencies, producing test cases that are syntactically invalid yet still encode meaningful execution intent. Rather than relying on one-shot generation to produce high-quality inputs, FuzzySQL treats test case construction as a refinement process and introduces a progressive error recovery pipeline. By repairing and completing partially valid SQL programs, the system systematically elevates test case quality and enables exploration of deeper, feature-dependent execution paths that are difficult to reach through direct generation alone. Embracing imperfection in exchange for broader exploration, FuzzySQL aligns naturally with the fundamental philosophy of fuzzing.
Table 1. SQL Error Handling Strategies
Error Category Definition Repair Strategy SAF RBR SAR
Duplicate Definition A data object is defined more than once.
Unsupported Feature Uses a feature or syntax not supported by the system.
Plugin/Component Errors Errors caused by failure or malfunction of plugins or components. Reload, reconfigure, or replace the plugin/component.
Inappropriate Setting A configuration setting is invalid or contextually incorrect. Adjust to a valid or recommended setting.
Formattable Errors Errors that can be resolved through formatting adjustments. Apply predefined repair rules through SQL statement formatting analysis.
Invalid Object Reference References a non-existent or illegal data object. Generates a SQL statement template to create or define the object and inserts it in front.
Preconditions Missing Required preconditions for execution are not met. Extract prerequisite requirements and build guiding prompts.
Incorrect Feature Usage Incorrect or improper usage of a feature. Attach the error messages directly to the corresponding SQL statements.
Violate Constraints Operation violates database constraints. Attach the error messages directly to the corresponding SQL statements.
SAF: Syntax-Aware Filtering RBR: Rule-Based Repair SAR: Semantic-Aware Repair
3.3.1. Error Taxonomy
To ensure the soundness of our repair strategy, we conduct a quantitative analysis of error messages returned by MySQL, a representative and widely deployed DBMS. From 12 hours of grammar-guided SQL generation, we collect 65,591 errors spanning 184 distinct types (out of 1,960 documented MySQL client error codes). Syntax-related errors account for 29.61% of all messages, while the top 50 semantic error types contribute 68.44%. These high-frequency semantic errors are manually analyzed and mapped to dedicated rule-based or intelligent semantic-aware repair strategies, while the remaining long-tail categories are handled via LLM-based semantic repair.
Following the same methodology, we perform 12-hour fuzzing sessions for other DBMSs and map the observed error patterns to the corresponding repair strategies. Although MySQL serves as the primary guide for designing our framework, FuzzySQL maintains compatibility across systems by adapting repairs based on error semantics rather than relying on DBMS-specific rules. The SQL error handling strategies we summarized are shown in Table 1.
3.3.2. Syntax-Aware Filtering
Before invoking any repair strategy, FuzzySQL performs error tagging to map DBMS error messages back to the corresponding SQL statements. This is achieved through log analysis and keyword alignment, allowing precise localization of faulty fragments for subsequent filtering or repair.
For syntax-related errors, FuzzySQL applies structural filtering to prioritize repair-worthy candidates. Specifically, SELECT statements with abstract syntax tree (AST) depth fewer than three are discarded as overly simplistic and unlikely to exercise deep execution logic. In contrast, statements involving DBMS-specific features (e.g., GTID, PROCEDURE, INSTALL COMPONENT) are always retained, as these often reside in under-tested paths and are more likely to expose non-trivial bugs. In addition, FuzzySQL filters out errors which are caused by duplicate object definitions or unsupported features rather than input errors. These include errors like Table 'XX'already exists, or This version of [DBMS] does not yet support 'XXX'. Filtering these cases early allows FuzzySQL to focus computational resources on repairing inputs with genuine structural or semantic promise.
3.3.3. Rule-Based Repair.
To efficiently handle frequent and structurally shallow error cases, FuzzySQL incorporates a rule-based repair layer that operates without requiring deep semantic modeling. This module targets predictable error patterns that can be corrected through localized adjustments—such as replacing invalid literals, reformatting malformed clauses, or adjusting configuration parameters—based solely on pattern matching and heuristic rewriting. Compared to semantic reasoning approaches, rule-based repair is lightweight and DBMS-agnostic. For example, errors such as ’Cannot get geometry object from data you send to the GEOMETRY field’ can be resolved by adjusting the literal assignment to conform to the expected data type. It avoids the overhead of parsing syntax trees or inferring dependencies, making it portable across dialects. More importantly, it enables fast recovery from common failures without invoking LLMs, thereby conserving LLM query budget for more complex, high-value repairs. By pre-staging repairable cases into this rule-based stage, FuzzySQL ensures high-throughput processing of grammar-generated inputs.
3.3.4. Intelligent Semantic-Aware Repair
For complex test failures rooted in semantic inconsistencies, such as undefined objects, unmet preconditions, or illegal feature combinations—FuzzySQL invokes a LLM to perform intelligent and context-sensitive repair. Unlike deterministic fixes, this stage leverages LLMs’ ability to interpret program intent and synthesize structurally valid SQL fragments to recover test cases that would otherwise be discarded.
The repair process begins by grouping remaining errors by type (e.g., invalid object reference, missing prerequisite, constraint violation), each associated with its corresponding faulty SQL fragment and execution context. For each case, FuzzySQL constructs a structured repair prompt that includes the SQL snippet, relevant schema history, the DBMS error message, and a natural language instruction describing the repair objective. Based on this prompt, the LLM generates a patch, either by inserting missing definitions (e.g., creating tables or functions) or rewriting invalid clauses. These transformations are injected directly into the test case for re-evaluation. For instance, when encountering an error such as PROCEDURE xxx does not exist, FuzzySQL constructs a prompt that instructs the LLM to synthesize a valid CREATE PROCEDURE template, placing it before the failing invocation. This effectively guides the repair process toward restoring missing database objects and leads to the generation of an interesting PROCEDURE.
In particular, LLM repairs are highly effective at recovering DBMS-specific constructs and operation sequences. This error-repair loop forms an iterative Chain-of-Thought (COT), where the model refines test cases based on error messages and repair strategies. This final repair phase enables FuzzySQL to recover high-value test inputs that exhibit deep structural complexity, significantly increasing the coverage of under-tested DBMS behaviors.
3.4. Replay-Guided Crash Validation
To ensure that detected crashes are genuinely reproducible and rooted in state-dependent behaviors of the DBMS, FuzzySQL implements a replay-guided crash validation pipeline followed by test case reduction. This mechanism isolates the root cause of failures and extracts a minimal yet effective proof-of-concept (PoC) that reproduces the crash.
Replay-based crash validation. Unlike traditional fuzzers that fork a new process for each test case, FuzzySQL resets the database state using DROP DATABASE and CREATE DATABASE statements, preserving the global session context across executions. While this approach improves performance, it introduces the possibility of residual state effects. In many cases, a crash does not occur when the current test case is executed in isolation, but only when certain prior statements have altered the environment. To address this, FuzzySQL first attempts to re-execute the crashing test case alone. If the crash does not reproduce, it replays the full execution history, executing prior statements in order before the crashing input, to verify whether the failure is state-dependent. If the error only manifests during the full replay, the issue is classified as a state-dependent crash and proceeds to minimization.
Dependency-preserving Test Case Reduction. Once a crash is confirmed, FuzzySQL automatically performs test case reduction to produce a minimal proof-of-concept (PoC). This process targets two dimensions: (1) structural simplification of individual SQL statements, and (2) reduction of the overall statement sequence. These two simplification strategies are applied iteratively and alternately, progressively minimizing the test case while ensuring the crash remains reproducible.
Reduction begins from the last statement in the sequence—where the crash is typically triggered—and proceeds backward. For structural simplification, each SQL statement is first parsed into an abstract syntax tree (AST), and then recursively simplified by pruning subtrees or nodes. The simplified AST is converted back into SQL and tested. If the crash still occurs, the simplification is retained. Otherwise, the system rolls back the change and attempts alternative simplifications. Once structural simplification weakens or removes certain data or control dependencies, the surrounding context becomes amenable to further reduction. At this stage, FuzzySQL attempts to delete preceding statements using a delta debugging strategy, preserving only those necessary for triggering the fault. The final output is a compact, semantically valid SQL sequence that isolates the root cause of the crash.
FuzzySQL’s replay and reduction workflow reconstructs the full crash path and eliminates irrelevant logic, producing a concise PoC of critical SQL statements, drastically improving debugging clarity and bug report quality.
4. Implemention
4.1. Prompt Engineering
To ensure stable generation quality and precise control over the output, FuzzySQL uses structured prompt templates tailored to SQL statement instantiation and semantic error repair. The complete prompt templates used in our implementation are provided in Appendix A.
4.1.1. Prompt Structure for SQL Instantiation
The goal of SQL instantiation is to transform grammar-expanded abstract templates into concrete, executable SQL queries. FuzzySQL constructs prompts with the following four key components:
  • Task Instruction: A concise natural language directive that instructs the LLM to generate a syntactically valid and semantically meaningful SQL query.
  • Schema Initialization Statements: Randomly generated CREATE TABLE/VIEW statements that define the database context. These provide schema and data type grounding (e.g., table names, column types, constraints) for coherent query generation.
  • SQL Templates: A set of SQL templates derived from the grammar, serving as a structural constraint.
  • Output Formatting Instruction: A final constraint that directs the LLM to return only complete SQL statements, avoiding comments, explanations, or multiple alternatives.
This structured prompting allows the LLM to fill in semantically appropriate content, such as meaningful column selections, valid WHERE clauses, or realistic JOIN clauses, all consistent with the provided schema.
4.1.2. Prompt Structure for Error Repair
If a SQL test case fails with errors addressable by semantic-aware repair, FuzzySQL constructs an LLM repair-instruction prompt containing:
  • Repair Instruction: A natural language request that instructs the model to identify the faulty SQL segment and either modify it or insert the required definitions to ensure successful execution.
  • Full SQL Context: The complete test case, including schema setup and execution sequence, providing the necessary context for understanding the failure and synthesizing a valid patch.
  • Error Annotations: The error message returned by the DBMS is injected near the offending SQL line, using inline comments.
  • Output Formatting Instruction: Follow the same formatting rules as those applied during the Instantiation phase.
The prompt structure allows the LLM to reason about schema-object relationships, identify the cause of failure, and propose actionable fixes, such as defining a missing object or correcting column usage, without discarding the broader intent of the test case.
4.2. Execution Framework and Instrumentation
FuzzySQL supports fuzzing for both client-server and embedded DBMS architectures through adaptive execution strategies tailored to the target under test. Instead of conventional stdin-based fuzzing, it delivers SQL inputs via realistic interaction paths to preserve semantic context and execution fidelity.
4.2.1. DBMS Fuzzing Drivers
FuzzySQL supports both client-server and embedded DBMS architectures through tailored execution strategies. For client-server systems (e.g., MySQL, MariaDB), FuzzySQL launches the command-line interface client via pexpect.spawn() and maintains a persistent session. To isolate test cases without restarting the server, it resets the environment using a standard SQL prelude: DROP DATABASE IF EXISTS test_db; CREATE DATABASE test_db; USE test_db;. This ensures a clean database context while preserving session-level state, enabling discovery of state-dependent bugs beyond fork-per-input fuzzers. For embedded engines (e.g., SQLite), FuzzySQL uses a forkserver-style subprocess to run the SQLite shell. It alternates between temporary and active database files (via .open tmp.db and .open test.db) to simulate a clean environment with minimal overhead, enabling efficient batch fuzzing without shell restarts.
4.2.2. Instrumentation and Coverage Feedback
All DBMS binaries are compiled with AFL++’s afl-clang-fast or afl-clang-fast++ toolchain to enable edge coverage instrumentation. AddressSanitizer (ASan) is also enabled to capture memory-related faults such as buffer overflows or use-after-free errors. No manual modifications to DBMS source code are required, allowing compatibility with standard upstream builds. Coverage feedback from AFL++ guides the mutation engine by prioritizing test cases that explore novel code paths. Only inputs that contribute new coverage are retained, while others are discarded.
4.2.3. Fuzzing Strategy and Determinism
FuzzySQL executes test cases in a strictly single-threaded fashion to maintain determinism, which is essential for reliable crash replay and debugging. Unlike AFL-style forkserver fuzzing, FuzzySQL maintains a persistent DBMS client and executes test inputs sequentially. This design ensures that the entire execution history is preserved and can be replayed when a crash occurs. In contrast, multi-threading is employed only for interacting with the LLM backend, where multiple concurrent threads handle prompt generation, SQL instantiation, and error repair. This architectural separation allows FuzzySQL to balance deterministic fuzzing with high-throughput generation. Crash detection relies on monitoring signal-based terminations (e.g., SIGSEGV, SIGABRT), abnormal exit codes, and standard error logs. Inputs that trigger these crashes, along with those that increase coverage, are retained in the seed pool for further mutation and replay. By combining persistent-mode execution, lightweight state resets, and compiler-based instrumentation, FuzzySQL achieves efficient and scalable fuzzing across diverse DBMS environments.
Figure 3. The number of new edges discovered in 24 hours.
5. Evaluation
5.1. Experimental Setup
All experiments are carried out on a server equipped with an Intel Xeon Gold 6430 CPU (32 cores), 256 GB RAM, and a NVIDIA A100 GPU, running Ubuntu 20.04 LTS. The GPU resources are reserved exclusively for local inference of LLMs. To ensure broad applicability, we evaluate FuzzySQL on five representative open source database systems that collectively cover both client-server and embedded DBMS architectures. Specifically, we target MySQL 9.1.0, MariaDB 11.4.0, SQLite 3.47.2, PostgreSQL 17.2, and ClickHouse v24.12.2.29-stable. FuzzySQL integrates the LLM model via a locally deployed instance of Qwen3-30B, running on the vLLM inference framework. During vulnerability mining, we also tested several advanced models; however, for controlled and reproducible evaluation, all reported results are based on Qwen3-30B. We used a temperature of 0.3–0.5 to balance stability and diversity. In pilot tests, a temperature of 0.4 produced natural yet varied outputs—sufficiently diverse for fuzzing under structured constraints. All experiments used an 8K-token context window.
To demonstrate the advantage of FuzzySQL, we compare it against three state-of-the-art open-source fuzzing baselines: Squirrel (Zhong et al., 2020), EET (Jiang and Su, 2024), and SQLancer (Rigger and Su, 2020a). Among them, Squirrel represents mutation-based fuzzing, while both EET and SQLancer adopt generation-based approaches. Since EET and SQLancer are not inherently a grey-box fuzzing tool, we first collect their generated test cases and then utilize FuzzySQL’s replay mechanism to measure its coverage. For SQLancer, we use TLP (Rigger and Su, 2020b) as the test oracle for MySQL, and NoREC (Rigger and Su, 2020a) for the other DBMS targets. This choice is constrained by the design of SQLancer, which employs different oracle-based testing strategies tailored to specific DBMSs. All baselines were evaluated under identical runtime conditions and DBMS versions. We adopted official configurations and standardized crash triage procedures to ensure fair and reproducible comparisons. Except for Squirrel, a mutation-based fuzzer using its default seeds, all baselines ran without a seed corpus. Each experiment was repeated five times on identical hardware, and the reported results are the averages, which show consistent statistical stability.
Table 2. Bugs discovered by FuzzySQL
DBMS Found Confirmed Fixed Feature-Related
MySQL 15 15 10 4
MariaDB 12 12 3 2
SQLite 2 2 1 1
ClickHouse 8 0 0 0
Total 37 29 14 7
5.2. Bug Discovery Effectiveness
To evaluate the practical effectiveness of FuzzySQL, we conducted a systematic fuzzing campaign on five representative open-source DBMSs. In total, FuzzySQL discovered 37 distinct bugs, among which 29 have been confirmed by vendors and 9 have been assigned CVE identifiers. As summarized in Table 2, 14 issues have already been fixed, while the remaining ones are scheduled for upcoming releases. Notably, 7 of the confirmed bugs are related to under-tested DBMS special features, such as GTID, PROCEDURE, and INSTALL COMPONENT, demonstrating FuzzySQL’s ability to explore feature-specific execution paths that are rarely exercised by existing fuzzers.
Beyond long-term bug discovery, we further compare fuzzing effectiveness under a fixed 24-hour budget. As shown in Figure 3, FuzzySQL consistently outperforms all baselines across all evaluated DBMSs in terms of newly discovered execution edges. The performance gap is particularly evident on MySQL, MariaDB, and PostgreSQL, where FuzzySQL achieves several times higher coverage than the strongest baseline. On SQLite and ClickHouse, the advantage is less pronounced, which can be attributed to the lightweight feature set of SQLite and the highly optimized execution engine of ClickHouse, where EET benefits from generating deeply nested analytical queries. Nevertheless, FuzzySQL still achieves the highest overall coverage across all systems.
To examine how this coverage advantage translates into concrete vulnerability discovery, Table 3 presents a breakdown of unique MySQL bugs discovered within 24 hours. Among the evaluated baselines, only EET uncovers a limited number of bugs, while none of the baselines discover any feature-related vulnerabilities. In contrast, FuzzySQL consistently exposes both feature-related and non-feature-related bugs within the same time budget. This result highlights a fundamental capability gap between FuzzySQL and existing fuzzers: while prior approaches can occasionally detect generic crashes, they lack the ability to synthesize and evolve stateful, feature-specific SQL programs required to exercise under-tested DBMS functionality.
Table 3. Unique MySQL bugs discovered by different fuzzers within 24 hours
[…]
FuzzySQL
Squirrel
EET
SQLancer
[…]
MySQL 3 $\circ$ Use After Free
MySQL 6 $\circ$ Assertion Failure
MySQL 8 $\bullet$ Assertion Failure
MySQL 9 $\circ$ NULL Ptr Deref
MySQL 13 $\bullet$ Assertion Failure
MySQL 15 $\circ$ NULL Ptr Deref
  • FR: $\bullet$ feature-related, $\circ$ not feature-related; !r: repair disabled; !m: mutation disabled.
  • $\dagger$ Bug ID corresponds to Table 4 (Appendix B).
Overall, these results demonstrate that FuzzySQL’s effectiveness lies not merely in discovering more bugs or achieving higher coverage, but in exposing qualitatively different classes of vulnerabilities that current DBMS fuzzers are structurally incapable of revealing. This capability stems from FuzzySQL’s integrated design, which enables efficient and systematic exploration of semantically sensitive logic across diverse database systems.
5.3. Ablation Study
We perform an ablation study to assess the impact of three core design choices in FuzzySQL: logic-shifting progressive mutation, automated semantic repair, and the choice of LLM for semantic instantiation. Our analysis addresses the following questions:
  • How critical is the logic mutation for triggering complex bugs?
  • Does the error repair pipeline meaningfully improve fuzzing effectiveness despite its computational cost?
  • How significantly does the choice of LLM affect the quality of generated test cases?
[…]
Figure 4. Ablation Study of FuzzySQL Components.
This observation aligns with Table 3. Under the same 24-hour budget, only the full FuzzySQL reliably discovers feature-related bugs and uncovers more unique vulnerabilities overall, while ablated variants miss several feature-dependent failures or exhibit reduced bug-finding capability. We attribute this difference to FuzzySQL’s integrated pipeline design, in which grammar-guided generation, progressive logic-shifting mutation, and semantic-aware repair jointly construct, evolve, and recover diverse execution contexts. Together, these components enable systematic exploration of feature-specific and state-dependent control paths and convert them into reproducible vulnerabilities, beyond what any individual component can achieve in isolation.
Effectiveness of Logic Mutation. We analyze all 29 confirmed bugs discovered by FuzzySQL to identify those that depend on logic-shifting transformations. A bug is considered logic-sensitive if its PoC contains a condition (e.g., =, IN, JOIN) mutated from another semantically valid form and becomes non-triggering when reverted. Using this criterion, 20 out of 29 bugs (68.96%) are logic-sensitive, indicating that logic mutation is a major contributor to vulnerability discovery. Figure 5 summarizes the distribution of these bugs by transformation type. NULL-related rewrites account for the largest fraction (11 PoCs), followed by equality polarity changes (10 PoCs) and join rewrites (8 PoCs). Membership predicates (IN/NOT IN) and other transformations, including AND/OR rewrites, ordering reversals, and EXISTS manipulation, contribute additional unique triggers. Together, these results show that small, semantics-preserving logic shifts can effectively expose alternative execution paths that are rarely exercised by structure-preserving or generation-only fuzzers.
Figure 5. Impact of Logic-shifting on POC Effectiveness.
Effectiveness of Automated Repair. Although semantic-aware repair introduces additional overhead, it substantially improves fuzzing yield. In our evaluation, 27% of the discovered bugs originate from test cases that were initially invalid and would have been discarded without repair, including cases with missing objects or unmet semantic preconditions.
Consistent with the system-level results and Table 3, disabling repair reduces the number of unique bugs discovered and causes several feature-related failures to be missed. These results indicate that automated repair is essential for preserving and exploiting high-value test cases, rather than serving as a mere convenience for handling invalid inputs.
Figure 6. Effectiveness of Different LLMs.
Impact of LLM Choice. We conducted an ablation study on MySQL to assess the impact of different LLMs on FuzzySQL’s performance. We evaluated Gemma3-27B, Gemma2-27B, Gemma2-9B, Qwen2.5-32B, and Qwen3-30B, each running on an A100 GPU with a context length of 8192. Over a 12-hour period, we recorded the number of valid SQL statements generated (Figure 6a) and the number of new execution edges discovered (Figure 6b). Figure 6a shows that Gemma2-9B produces the largest number of valid SQL statements with the highest throughput and efficiency. However, Figure 6b reveals that this advantage does not translate into deeper code exploration. In contrast, Qwen3-30B generates fewer statements overall but consistently achieves the highest coverage growth over time. This indicates that its outputs are of higher semantic quality, capable of exercising deeper and more diverse program behaviors.
These results demonstrate that stronger models, such as Qwen3-30B, produce higher-quality and more diverse test cases, leading to greater coverage and bug-discovery potential. However, we opt for locally deployed models rather than API-based SoTA ones (e.g., GPT-4o) due to cost and scalability concerns—running such models at fuzzing scale would incur thousands of dollars and hundreds of compute hours. Similarly, reasoning-oriented models (e.g., OpenAI-o1, DeepSeek-R1) were excluded because their multi-step reasoning severely limits throughput, making them unsuitable for large-scale fuzzing. Therefore, FuzzySQL’s design favors efficient, instruction-aligned local LLMs, which strike a balance between semantic quality and generation efficiency for large-scale fuzzing workloads.
5.4. Case Study
To complement our quantitative evaluation, we present case studies that demonstrate the effectiveness of FuzzySQL in three distinct dimensions. These examples illustrate how the system exposes critical vulnerabilities in the DBMS internals by leveraging feature-specific generation, semantic-aware mutation, and error recovery. Each case highlights a different challenge in DBMS fuzzing and how FuzzySQL overcomes it through its design.
Feature-related bugs in under-tested DBMS functionalities. We present two representative vulnerability-triggering inputs found in MySQL and MariaDB, which are shown in Listing 3, Listing 4 and 5 respectively. These cases violate the standard SQL syntax or DBMS usage rules—MySQL does not support the combined use of DESC and FORMAT; MariaDB’s KILL statement lacks support for predicate logic; and the incorporation of composite subquery expressions within PROCEDURE declarations in MariaDB is improper. As a result, traditional DBMS fuzzers, often guided by hand-crafted grammars and strict semantic correctness, are unlikely to generate such inputs.
In contrast, FuzzySQL leverages grammar-driven structural expansion combined with LLM-based forced instantiation to generate atypical, boundary-pushing SQL inputs. By decoupling structure from content, these ”fuzzy templates” serve as scaffolds that guide the LLM beyond its default bias toward well-formed, canonical queries—a limitation inherited from pretraining on normative data. This mechanism encourages the generation of rule-bending or underspecified statements that lie outside conventional syntax but remain executable. Such flexibility cannot be easily achieved through manually crafted grammars. For example, the combination of KILL with predicate logic in Listing 4 represents a usage pattern that is both rare and counterintuitive, something human-authored templates are unlikely to anticipate, and traditional generators lack the means to instantiate. FuzzySQL’s design, by contrast, allows it to discover these error-prone interactions naturally through guided structural diversity and semantic inference.
1CREATE TABLE t1 (v1 JSON, v2 CHAR(7), v3 BINARY(3));
2DESC ANALYZE FORMAT = JSON (SELECT v1 FROM t1 UNION XXXX);
Listing 3: MySQL crash via invalid DESC ANALYZE syntax.
1KILL 1 IN (SELECT 1) IN (XXXX);
Listing 4: MariaDB crash via malformed KILL expression.
1CREATE PROCEDURE test_proc (id VARCHAR(255)) BEGIN DECLARE dt
2DATETIME(6) DEFAULT ROW(XXXX) = SOME(SELECT 1) = ALL(XXXX);
Listing 5: MariaDB crash via subquery misuse in PROCEDURE.
Logic-sensitive bugs enabled by conditional mutation. FuzzySQL’s lightweight logic mutation enables subtle yet effective transformations in SQL conditions, allowing it to explore alternative control paths often missed by traditional fuzzers. A representative example is shown in Listing 6, where a conditional expression inside an INSTALL COMPONENT statement includes a comparison of the disjunctive subquery. This particular crash is triggered only when the clause t0.c1 = 1 is present—replacing it with t0.c1 != 1 or any other variation prevents the assertion failure (as confirmed by GDB stack trace analysis). This demonstrates that the bug is not merely exposed through structural coverage, but depends on a precise logical condition being satisfied during execution. Existing mutation strategies, whether focused on AST structures or statement-level heuristics, often overlook such fine-grained control flow variations and struggle to achieve this level of logical precision. By systematically shifting operators, FuzzySQL introduces semantic perturbations that preserve syntax while altering the evaluation paths. This approach not only increases test diversity, but also surfaces fragile logic handling within DBMS internals.
Repaired bugs salvaged from initially invalid test cases. FuzzySQL’s semantic-aware repair module enables recovery of otherwise invalid inputs that can still uncover real vulnerabilities. Listing 7 shows a PoC that triggers a heap-buffer-overflow in SQLite during join resolution in sqlite3WhereEnd(). Initially, this input failed with the error “ambiguous column name: main.t2.v1” due to multiple unaliased uses of table t2, which caused semantic ambiguity during join resolution. Traditional fuzzers typically discard such malformed cases. Instead, FuzzySQL follows the SAR Violate Constraints error handling strategy to capture error message via tagging and incorporates it into a structured LLM prompt. The model identifies the ambiguity and suggests a minimal fix—assigning an alias to one instance of t2—which enables execution and exposes the underlying bug. This example highlights how FuzzySQL’s repair pipeline can salvage high-value test cases that would otherwise be lost, turning parse-time failures into exploitable paths through lightweight, context-aware correction.
6. Discussion
This section discusses the broader implications, strengths, and limitations of FuzzySQL. This work not only delivers concrete results in uncovering functionality-related vulnerabilities in DBMSs, but also offers new perspectives on how LLMs can reshape the design of fuzzing frameworks.
Uncovering Hidden Vulnerabilities in DBMS Special Features. Our findings reveal that modern DBMSs expose a wide range of special-purpose SQL features—such as replication control, logging, and statistical commands—that are rarely targeted by traditional fuzzers. Although infrequent in everyday workloads, these features are tightly integrated into critical control paths and can cause severe runtime failures when misused or exercised under uncommon conditions. FuzzySQL systematically explores this under-tested space by generating multi-statement SQL sequence that simulate realistic yet edge-case feature combinations. This exposes control-path inconsistencies and metadata mismanagement bugs that are not triggered by typical DML/DDL queries. These results highlight a critical blind spot in conventional fuzzing techniques, which tend to focus on common syntax patterns and fail to explore the operational depth of feature-rich systems.
Flexibly Adapt to Diverse DBMS Targets. To migrate FuzzySQL to a new DBMS, users only need to provide the target’s grammar specification (e.g., an ANTLR4 .g4 file), from which FuzzySQL automatically extracts syntax rules for template-based SQL generation. In addition, FuzzySQL requires a lightweight database initialization component to generate basic CREATE and INSERT statements. This component involves enumerating supported data types of the target DBMS and applying largely uniform schema and data generation rules, which introduces minimal manual effort and is straightforward to implement. For error repair, users collect execution feedback during an initial fuzzing phase and map observed errors to high-level categories. Most repair strategies are DBMS-agnostic and can be reused across systems, requiring little to no customization for new targets. As a result, supporting new DBMS targets requires limited manual customization.
1INSTALL COMPONENT ’plugin2’ SET GLOBAL default.variable1 = ON, t0.c1 = 1 OR 1 = ANY (XXXX);
Listing 6: MySQL crash via malformed INSTALL COMPONENT inducing conditional logic.
1CREATE TABLE t1 (XXXX);
2CREATE TABLE t2 (XXXX) WITHOUT ROWID;
3INSERT INTO t2 (XXXX) VALUES (XXXX);
4ANALYZE t2;
5WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 ON (ref_2.v2 = 1))); -- Error: ambiguous column name: main.t2.v1
6WITH cte_0 AS (SELECT 1) SELECT DISTINCT ref_2.v1 FROM (cte_0 JOIN generate_series(79, 105, 7) AS ref_1 ON (ref_1.value = 1)) CROSS JOIN (t2 AS t2a JOIN (t1 AS ref_2 JOIN t2 as ref_3 ON (ref_2.v2 = 1))); -- Fixed and Trigger Crash
Listing 7: Repaired SQLite input that originally failed due to ambiguous column references.
Revisiting the Role of LLMs in Fuzzing. A key insight of this work is that LLMs, when guided by structured grammar templates and diverse execution contexts, can serve as more than just language generators—they can be used as semantic reasoning engines. Rather than relying solely on random AST-level mutations, FuzzySQL uses LLMs to instantiate meaningful programs based on context, such as generating valid stored procedure bodies or logically connected statement sequences. This allows fuzzing to move beyond surface-level syntax exploration and into deeper logic modeling. By decoupling structural constraints (enforced by templates) from content generation (handled by LLM), we retain structural diversity while still harnessing the semantic strengths of LLMs. This design helps mitigate LLMs’ bias toward producing safe, canonical outputs and encourages exploration of more unusual and error-prone cases.
Limitations and Future Outlook. While FuzzySQL is effective in exposing semantic and control-flow bugs, it currently underperforms compared to EET in metrics such as statement depth and nesting complexity. This stems from inherent limitations in current LLMs, which tend to struggle with generating deeply nested or highly recursive SQL structures. Additionally, LLM-generated inputs occasionally require repair due to inconsistencies between declared objects and later usage, prompting our design of a repair pipeline. However, these challenges largely reflect the limitations of current LLMs. As LLMs improve in reasoning depth and code synthesis capability, their effectiveness in fuzzing high-complexity structures will improve correspondingly. FuzzySQL excels at triggering feature-related failures and hard-to-reach logic branches, aspects that are often invisible to basic coverage models.
7. Related Work
7.1. DBMS Fuzzing
Fuzzing has emerged as a critical technique for testing database management systems (DBMSs) by directly targeting components such as parsers, optimizers, and execution engines through crafted SQL inputs. Depending on the bug types targeted—ranging from crash vulnerabilities (37; X. Liu, X. Kong, L. Liu, and K. Chiang (2018); R. Zhong, Y. Chen, H. Hu, H. Zhang, W. Lee, and D. Wu (2020); M. Wang, Z. Wu, X. Xu, J. Liang, C. Zhou, H. Zhang, and Y. Jiang (2021); J. Fu, J. Liang, Z. Wu, M. Wang, and Y. Jiang (2022); Z. Jiang, J. Bai, and Z. Su (2023); J. Liang, Y. Chen, Z. Wu, J. Fu, M. Wang, Y. Jiang, X. Huang, T. Chen, J. Wang, and J. Li (2023); J. Fu, J. Liang, Z. Wu, and Y. Jiang (2024)) to logic inconsistencies (Rigger and Su, 2020c, b, a; Liang et al., 2022; Ba and Rigger, 2023; Jiang and Su, 2024; Zhang and Rigger, 2025) and performance regressions (Jung et al., 2019; Liu et al., 2022; Zhang et al., 2022; Ba and Rigger, 2024)—existing DBMS fuzzers follow either generation-based or mutation-based strategies.
Generation-based approaches rely on hand-crafted grammars or learned rules to synthesize SQL sequences from scratch. SQLSmith (37) uses randomized AST construction to generate structurally valid SQL inputs, but often fails on semantic soundness. To enhance expressiveness, DynSQL (Jiang et al., 2023) dynamically tracks schema evolution to guide generation, while TreeGAN (Liu et al., 2018) adopts grammar-aware GANs to synthesize syntactically valid SQL trees. SQLancer (Rigger and Su, 2020c, b, a; Zhang and Rigger, 2025) and EET (Jiang and Su, 2024) generate ASTs while preserving contextual variable bindings to ensure validity. However, most generators still rely heavily on manual rule engineering or specifications. The concurrent work ShQveL (Zhong and Rigger, 2025) uses LLM to construct fragments and then uses these fragments to combine SQL statements. This test case generation method still relies on the implementation of general syntax rules.
Mutation-based fuzzers, in contrast, transform existing SQL statements through changes at the syntax or intermediate representation (IR) level. Some prior work also explores context-dependent mutations arising from the composition of SQL statement sequences. Squirrel (Zhong et al., 2020) represents SQL as intermediate representation (IR) and applies typed, data-aware mutation. RATEL (Wang et al., 2021) adds fine-grained coverage tracking and feedback-guided deduplication for robust crash reporting. Griffin (Fu et al., 2022) shuffles SQL statement sequences while tracking metadata dependencies to enable semantically consistent recombination. LEGO (Liang et al., 2023) introduces a type affinity model to mutate SQL type sequences and enrich structural diversity, though semantic restoration remains limited. BUZZBEE (Yang et al., 2024) generalizes mutation by abstracting all database interactions into define–use–invalidate triples and uses a lightweight query language (CQL) to ensure semantic compatibility.
Unlike prior DBMS fuzzers that rely on rigid rule coding or handcrafted dependency models, FuzzySQL leverages LLMs to infer semantic dependencies and supports grammar-guided test case synthesis, targeting feature-rich, stateful DBMS.
7.2. Grammar-Based Fuzzing
Grammar-based fuzzing generates syntactically valid test inputs by leveraging formal grammars such as context-free grammars (CFGs), addressing the limitations of traditional mutation-based fuzzers that often fail in structured input formats. Early tools like LangFuzz (Holler et al., 2012), CSmith (Yang et al., 2011), NAUTILUS (Aschermann et al., 2019),and Superion (Wang et al., 2019) and FuzzFlow (Xu et al., 2024) established grammar-guided mutation or derivation-based input synthesis to maintain structural correctness. Later works incorporated context-sensitivity and probabilistic modeling to capture semantic constraints more effectively, such as Skyfire (Wang et al., 2017), SAGE (Zhou et al., 2023), and GrayC (Even-Mendoza et al., 2023), which learn constraints from corpora or feedback. Other systems like SYNTHFUZZ (Limpanukorn et al., 2024), FreeDom (Xu et al., 2020), and MLIRSmith (Wang et al., 2023) encode def-use chains, type consistency, or intermediate representations to preserve semantic validity.
Despite these advances, many grammar-based fuzzers still struggle with expressing inter-statement dependencies, environmental assumptions, and stateful execution paths—factors critical in domains like database fuzzing. Some systems mitigate this via probabilistic grammars (Zhang et al., 2024b), derivation tree splicing (Alsaeed and Young, 2023), or grammar automata (Srivastava and Payer, 2021), but semantic gaps remain. Tribble (Havrikov and Zeller, 2019) systematically explores derivation paths to maximize rule coverage. IR-based systems such as POLYGLOT (Chen et al., 2021), MLIRod (Suo et al., 2024), and NNSmith (Liu et al., 2023) aim to generalize semantics, while Grimoire (Blazytko et al., 2019), Evogram (Gopinath et al., 2021) and Grammar-based Whitebox Fuzzing(Godefroid et al., 2008) explore structure without formal grammars. However, integrating semantics into generation pipelines remains complex.
While traditional grammar-based fuzzers emphasize syntactic validity, FuzzySQL synthesizes diverse grammar-derived templates and then uses LLMs to force context-aware completion, enabling the generation of corner cases that uncover functionality-specific bugs beyond the expressive reach of fixed grammar rules.
7.3. LLM-Assisted Fuzzing
LLMs have recently enabled a new generation of fuzzing tools capable of generating semantically rich and structurally valid inputs across a wide range of domains. For example, CovRL-Fuzz (Eom et al., 2024) combines reinforcement learning with LLM-based mutation to improve coverage in JavaScript engines. PromptFuzz (Lyu et al., 2024) mutates prompts instead of programs to synthesize valid C/C++ fuzz drivers guided by runtime feedback. Fuzz4All (Xia et al., 2024) introduces automatic prompting to fuzz heterogeneous systems without manual seed or grammar engineering. In protocol fuzzing, LLMIF (Wang et al., 2024b), CHATAFL (Meng et al., 2024), and mGPTFuzz (Ma et al., 2024) extract protocol structure from natural language documents to guide specification-aware input generation. TitanFuzz (Deng et al., 2023), FuzzGPT (Deng et al., 2024), KernelGPT (Yang et al., 2025), and ProphetFuzz (Wang et al., 2024a) target deep learning frameworks and OS kernels by mining vulnerable code patterns or system configurations using LLM reasoning.
Additional efforts explore LLMs as intelligent mutators or input generators. MetaMut (Ou et al., 2024) synthesizes compiler-level mutators from AST APIs and domain prompts. Magneto (Zhou et al., 2024) leverages LLMs to reconstruct multi-stage call chains for exploit generation. ECG (Zhang et al., 2024a) uses LLMs to extract syscall specifications and guide semantic mutations in embedded OSs, while CodaMosa (Lemieux et al., 2023) mixes search-based testing with Codex-based test generation. EAGLEYE (Liu et al., 2025) targets hidden interfaces in IoT firmware by extracting routing patterns, and InputBlaster (Liu et al., 2024) simulates edge-case input behavior in mobile apps via constraint-violating prompt synthesis.
Some studies focus on improving seed quality and diversity (Oliinyk et al., 2024). Sedar (Fu et al., 2024) leverages LLMs for SQL seed adaptation across DBMS dialects but suffers from context hallucination and dependency misalignment.
Notably, FuzzySQL does not rely on LLM for mutation, but instead leverages them selectively for semantic instantiation and error-guided repair. This integration enables exploration of deeper feature-related paths that remain out of reach for most prompt-only or grammar-guided LLM fuzzers.
8. Conclusion
We presented FuzzySQL, an LLM-driven fuzzing framework that uncovers deep-seated bugs in DBMS special features often missed by existing tools. By combining grammar-guided generation, semantic mutation, hybrid error repair, and state-aware execution, FuzzySQL enables broad feature coverage and precise bug reproduction. Evaluations across multiple DBMSs revealed 37 unique vulnerabilities, 7 of which were tied to under-tested DBMS features. These results validate the effectiveness of our approach and highlight the need for semantically guided fuzzing in database security. Our findings suggest that integrating LLMs with structured generation offers a promising direction for advancing semantic fuzzing across complex software systems.
Appendix
Appendix A Prompts Used In FuzzySQL
This section documents the exact prompts used in FuzzySQL for SQL instantiation and semantic-aware repair.
SQL Instantiation Prompt Here we initialize a database test_db by executing the initialization SQL statements:
{init_schema_statements}
Please instantiate the following SQL statements using the guidance of the given SQL templates:
{sql_templates}
The generated SQL statements must satisfy the following requirements: They must be syntactically and semantically correct. They must be executable in {targetDB}. Each subsequent statement must reference only objects created by previous statements (e.g., tables, views, or columns). You are allowed to complete the necessary parts of the templates. Return only the SQL statements generated from the templates. Do not repeat the initialization SQL statements or include any other content. Before output, correct any syntax or semantic errors in the generated SQL statements. Output the result in JSON format. Example response:
["SQL1;", "SQL2;", ...]
Semantic-Aware Repair Prompt We need you to fix a SQL test case that contains erroneous SQL statements. The input test case contains multiple SQL statements. SQL statements that require repair are explicitly marked using the following format:
-- [Need to repair<
<SQL statement>
-- <error message>
-- (repair suggestion)
-- >Need to repair]
Here is the input test case:
{casecontent}
You should fix each erroneous SQL statement using: the surrounding SQL context, the error message, and the repair suggestion if provided. Target database: {targetDB}. Do not output anything other than SQL statements. Output the fully fixed test case in JSON format. Example response:
["SQL1;", "SQL2;", ...]
Appendix B Vulnerability Status Across DBMSs
This section summarizes the vulnerabilities discovered across multiple DBMSs, and reports their affected components and current handling status (e.g., fixed, verified, or in progress). It provides an at-a-glance view of our disclosure and validation outcomes.
Table 4. Vulnerability Discovered Across DBMSs
DBMS ID Reference Location Status
MySQL 1 CVE-redacted Optimizer Fixed
2 CVE-redacted Optimizer Fixed
3 CVE-redacted Optimizer Fixed
4 CVE-redacted Optimizer Fixed
5 CVE-redacted Optimizer Fixed
6 CVE-redacted Optimizer Fixed
7 CVE-redacted Optimizer Fixed
8 CVE-redacted Components Services Fixed
9 CVE-redacted Optimizer Fixed
10 S{redacted} Item_ref Fixed
11 S{redacted} Tmp_table Verified
12 Debug Build MDL_checker Verified
13 Debug Build Mysql_execute_command Verified
14 Debug Build Explain_query Verified
15 Debug Build Item_in_subselect Verified
SQLite 1 {redacted} WhereEnd Fixed
2 {redacted} VdbeExec Fixed
MariaDB 1 MDEV-redacted Optimizer Fixed
2 MDEV-redacted Tmp_table Fixed
3 MDEV-redacted Item_func_not Fixed
4 MDEV-redacted Item_subselect Verified
5 MDEV-redacted Item_subselect Verified
6 MDEV-redacted Item Verified
7 MDEV-redacted Item_field Verified
8 MDEV-redacted Filesort Verified
9 MDEV-redacted Select_value_catcher Verified
10 MDEV-redacted Window_funcs_sort Verified
11 MDEV-redacted Optimizer Verified
12 MDEV-redacted Filesort Verified
ClickHouse 1 {redacted} Condition_variable In progress
2 {redacted} FunctionIn In progress
3 {redacted} FunctionBinaryArithmetic In progress
4 {redacted} FormatStringHelperImpl In progress
5 {redacted} AlterCommand In progress
6 {redacted} ExecuteQueryImpl In progress
7 {redacted} Planner In progress
8 {redacted} IAST In progress
Status legend. We label an entry as Fixed once a patch is available in an upstream release; Verified indicates that we reproduced the issue and reported it, but an official fix is still pending; and In progress denotes that triage with the maintainers is ongoing. Overall, Table 4 summarizes 37 issues across four DBMSs (MySQL: 15, SQLite: 2, MariaDB: 12, ClickHouse: 8), with identifiers redacted where required by responsible disclosure.
Appendix C Open Science
This section follows the ACM CCS Open Science policy by listing the artifacts needed to evaluate our core contributions and explaining how reviewers can access them during double-blind review.
Access during double-blind review.
We provide an anonymized artifact repository at:
https://anonymous.4open.science/r/fuzzysql-D589/
Artifacts.
The repository contains the artifacts needed to evaluate this paper, including: (i) the implementation of FuzzySQL, (ii) experiment scripts/configurations, (iii) instructions to run the experiments.
Unavailable artifacts and responsible disclosure.
This work studies real-world security vulnerabilities. Detailed information about specific vulnerabilities (e.g., vulnerability-triggering inputs, proof-of-concept exploits, and affected systems) is not publicly disclosed at submission time due to responsible disclosure and security concerns. Several of the discovered vulnerabilities have been reported to the corresponding vendors and are undergoing coordinated disclosure. Releasing full vulnerability details prior to patch availability could pose security risks.
To enable evaluation of our methodology, we provide the full implementation of FuzzySQL, experiment scripts, configurations, and partially sanitized examples that illustrate the vulnerability patterns discussed in the paper, without exposing sensitive or actionable details.